OpenStudy (anonymous):

Which of the following is a correct equation for the line passing through the point (-3,2) and having slope m = 2/3? Check all that apply. A. 2x-3y=-12 B. y-2=2/3(x+3) C. y=2/3x D. y=2/3x+4

OpenStudy (anonymous):

use point slope formula y-y1=m(x-x1)

OpenStudy (anonymous):

please help:/

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

y-2=2/3(x--3) y-2=2/3x+2 y=2/3x+4 thus D

OpenStudy (anonymous):

but isn't there more than one answer? because on the question it says check all that apply?

OpenStudy (anonymous):

yeah A,B and D A is in standard form B is in point slope form D is in slope intercept form

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no problem